BIMA ARRAY MOLECULAR SEARCHES IN COMET HALE-BOPP(C/1995 O1)

L. E. SNYDER, \underline{A. MEYERTHOLEN}, J. M. VEAL, Department of Astronomy, University of Illinois, Urbana IL 61801; Y.-J. KUAN, Department of Earth Sciences, National Taiwan Normal University, and IAA, Taipei, Taiwan; I. DE PATER, J. R. FORSTER, M. C. H. WRIGHT, Department of Astronomy, University of California, Berkeley, CA 94720; L. M. WOODNEY, M. F. A’HEARN, Department of Astronomy, University of Maryland, College Park MD 20742-2421; and P. PALMER, Department of Astronomy and Astrophysics, University of Chicago.

Cometary molecules specifically targeted for observation with the BIMA Array during the 1997 apparition of comet Hale-Bopp included HCN, HCO^{+}, and CS, but because the BIMA Array has a very flexible, wideband spectrometer^{a}, many other species had transitions in the bandpass which could be observed simultaneously with the target molecules. We discuss the results of these “search” observations which included transitions of AlCl, CO, ^{13}CS, ^{34}SO, SiO, ^{29}SiS, C_{2}H, C_{2}S, HCS^{+}, H^{15}NC, MgNC, OCS, SO_{2}, ^{33}SO_{2}, ^{34}SO_{2}, Si^{13}CC; C_{3}H, C_{3}S, HOCO^{+}, C_{3}H_{2}, C_{4}H, HC_{3}N, C_{5}H, CH_{3}CN, CH_{3}^{13}CN, CH_{3}OH, NH_{2}CHO, C_{6}H, CH_{2}CHCN, HCOCH_{3}, HCOOCH_{3}, (CH_{3})_{2}O, CH_{3}CH_{2}OH, and HC_{7}N (here, a molecular formula with no mass number denotes the main isotopomer).

This work was partially funded by: NASA NAG5-4292, NAG5-4080, and NGT5-0083; NSF AST 96-13998, AST96-13999, AST96-13716, and AST96-15608; Taiwanese grants NSC 86-2112-M-003-T and 87-2112-M-003-007; and the Universities of Illinois, Maryland, and California, Berkeley.

aW. J. Welch et. al., PASP, 108, 93 (1996)